In each reaction box, place the best reagent and cneed help with the reagents and conditions!
Answer #1
Concepts and reason

The problem is based on the concept of elimination and substitution of bromine.

This can be done by taking use of appropriate reagents.

Fundamentals

For brominating, Br2{\rm{B}}{{\rm{r}}_2} is used, the equivalent of Br2{\rm{B}}{{\rm{r}}_2} used depends on the demand of product.

Sodium amide, NaNH2{\rm{NaN}}{{\rm{H}}_{\rm{2}}} is strong base used for abstraction of proton.

Part a

The reaction is as follows:

Excess NaNH2
Br

The reaction is as follows:

e o
Na
leq Br2
Br
Br

Part b

The reaction is as follows:

H-C=C-H-
NaNH
?
>H-C = CºNaⓇ

The reaction is as follows:

H-C= CºNaº_CH3CH,Br
>H-C=C-CH2CH3

The reaction is as follows:

Br
H-C=C-H-CH-CH-
HBr,2eq
H.-C-C-CH,CH;

Part c

The reaction is as follows:

The reaction is as follows:

© ©
-C=C Na
excess NaNH,
Br

The reaction is as follows:

_C=C Na
K
C
=C-CH,
CH,Br

The reaction is as follows:

C=C-CH,
oro, o
| Li, NH

Ans: Part a

The complete reaction with suitable reagents is as follows:

1)Excess NaNH2
2)leq Br2
Br
Br
Br

Part b

The complete reaction with suitable reagents is as follows:

1)NaNH,
2)CH2CH Br
3 )HBr, 2–eq
H-C=C-H-
Br
Br

Part c

The complete reaction with suitable reagents is as follows:

1)Br2
2 )Excess NaNH,
3)CH Br
4)Li, NH3(1)